# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ 2 Preliminaries
By #$\#S$ we will denote the cardinality of a finite set $S$. The finite field of two elements will be denoted by $\mathbb{F}_{2}$, and for a positive integer $n$, $\mathbb{F}_{2}^{n}$ will denote the vector space of dimension $n$ over $\mathbb{F}_{2}$. By $\oplus$, we will denote the addition operator over both $\mathbb{F}_{2}$ and $\mathbb{F}_{2}^{n}$. An element of $\mathbb{F}_{2}^{n}$ will be considered to be an $n$-bit binary string.
For $n\geq 0$, let $\mathbf{x}=(x_{1},\ldots,x_{n})$ be an $n$-bit binary string. The support of $\mathbf{x}$ is ${\sf supp}(\mathbf{x})=\{1\leq i\leq n:x_{i}=1\}$, and the weight of $\mathbf{x}$ is ${\sf wt}(\mathbf{x})=\#{\sf supp}({\sf supp}(\mathbf{x}))$. By $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$ we will denote the all-zero and all-one strings of length $n$ respectively. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ be two $n$-bit strings. The distance between $\mathbf{x}$ and $\mathbf{y}$ is $d(\mathbf{x},\mathbf{y})=\#\{i:x_{i}\neq y_{i}\}$; the inner product of $\mathbf{x}$ and $\mathbf{y}$ is $\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}\oplus\cdots\oplus x_{n}y_{n}$; and we define $\mathbf{x}\leq\mathbf{y}$ if $x_{i}\leq y_{i}$ for $i=1,\ldots,n$.
@@
Let $f$ be an $n$-variable function. The algebraic normal form (ANF) representation of $f$ is the following: $f(X_{1},\ldots,X_{n})=\bigoplus_{\bm{\alpha}\in\mathbb{F}_{2}^{n}}a_{\bm{\alpha}}\mathbf{X}^{\bm{\alpha}}$, where $\mathbf{X}=(X_{1},\ldots,X_{n})$; for $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}_{2}^{n}$, $\mathbf{X}^{\bm{\alpha}}$ denotes the monomial $X_{1}^{\alpha_{1}}\cdots X_{n}^{\alpha_{n}}$; and $a_{\bm{\alpha}}\in\mathbb{F}_{2}$. The (algebraic) degree of $f$ is ${\sf deg}(f)=\max\{{\sf wt}(\bm{\alpha}):a_{\bm{\alpha}}=1\}$. Functions of degree at most 1 are said to be affine functions. Affine functions with $a_{\mathbf{0}_{n}}=0$ are said to be linear functions. It is known that if $f$ is balanced, then ${\sf deg}(f)\leq n-1$.
@@
For two $n$-variable functions $f$ and $g$, the distance between them is $d(f,g)=\#\{\mathbf{x}\in\mathbb{F}_{2}^{n}:f(\mathbf{x})\neq g(\mathbf{x})\}$. The nonlinearity of an $n$-variable function $f$ is ${\sf nl}(f)=\min d(f,g)$, where the minimum is over all $n$-variable affine functions $g$.
@@
A function is said to be plateaued if its Walsh transform takes only the values $0,\pm v$, for non-zero $v$. We define the linear bias of an $n$-variable Boolean function $f$ to be ${\sf LB}(f)=1/2-{\sf nl}(f)/2^{n}$.
@@
Let $f$ be an $n$-variable function. The algebraic immunity of $f$ is defined [23, 44] as follows: ${\sf AI}(f)=\min_{g\neq 0}\{{\sf deg}(g):\mbox{ either }gf=0,\mbox{ or }g(f\oplus 1)=0\}$. It is known [23] that ${\sf AI}(f)\leq\lceil n/2\rceil$.
The fast algebraic attack (FAA) was introduced in [22]. The idea of the attack is based on the following observation. Let $f$ be an $n$-variable function and suppose $g$ is another $n$-variable function of degree $e$ such that $gf$ has degree $d$. If both $e$ and $d$ are small, then $f$ is susceptible to an FAA. Given $f$, and for $e$ and $d$ satisfying $e+d\geq n$, it is known [22] that there exists functions $g$ and $h$ with ${\sf deg}(g)=e$ and ${\sf deg}(h)\leq d$ such that $gf=h$. For any pair of functions $g$ and $h$ of degrees $e$ and $d$ respectively, satisfying $gf=h$, we have $h=gf=gf^{2}=(gf)f=hf$ and so if $h\neq 0$, then $(1\oplus h)$ is an annihilator of $f$ which implies $d\geq{\sf AI}(f)$. If $1\leq e<{\sf AI}(f)$, then $g$ is not an annihilator of $f$ and so $h=gf\neq 0$, and we obtain $e+d\geq 1+{\sf AI}(f)$. The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\allowbreak\{{\sf deg}(g)\allowbreak+\allowbreak{\sf deg}(fg):\allowbreak 1\allowbreak\leq\allowbreak{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$. For any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
${\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$ $=$ $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})=\pi_{1}(\mathbf{X})Y_{1}\oplus\cdots\oplus\pi_{m}(\mathbf{X})Y_{m}\oplus h(\mathbf{X}).$ (2)
@@
${\sf AI}({\sf MM}_{2m})$$\geq$ ${\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}.$
@@
00 0 $=$ $g(\mathbf{X},\mathbf{Y}){\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$ (3)
@@
Since the right hand side of (3) is equal to 0, for $\bm{\omega}\in\mathbb{F}_{2}^{m}$, the coefficient of $\mathbf{Y}^{\bm{\omega}}$ in the expansion on the right hand side of (3) must be equal to 0. Since $g(\mathbf{X},\mathbf{Y})\neq 0$, let $w\geq 0$ be the minimum integer such that there is an $\bm{\omega}^{\star}$ with ${\sf wt}(\bm{\omega}^{\star})=w$ and $g_{\bm{\omega}^{\star}}(\mathbf{X})\neq 0$. In (3), equating the coefficient of $\mathbf{Y}^{\bm{\omega}^{\star}}$ to 0, we have
00 0 $=$ $g_{\bm{\omega}^{\star}}(\mathbf{X})\left(h(\mathbf{X})\oplus\left(\bigoplus_{i\in{\sf supp}(\bm{\omega}^{\star}))}\pi_{i}(\mathbf{X})\right)\right)$
$=$ $g_{\bm{\omega}^{\star}}(\mathbf{X})\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}.$
So $g_{\bm{\omega}^{\star}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})$. Consequently, ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\star})+{\sf deg}(g_{\bm{\omega}^{\star}})\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}$.
If, on the other hand, $g(\mathbf{X},\mathbf{Y})$ is an annihilator for $1\oplus{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$, then a similar argument shows that $g_{\bm{\omega}^{\star}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus 1\oplus h(\mathbf{X})$, and again we have ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big{)}$.
@@
We consider a very well known extension of ${\sf MM}_{2m}$ to odd number of variables.
@@
The first point is well known. The second point is immediate from the definition of ${\sf MM}_{2m+1}$. We provide the proof of the third point. For brevity, let us write $\mathbf{Z}=(\mathbf{X},\mathbf{Y})$. Clearly if $g(\mathbf{Z})$ is an annihilator for ${\sf MM}_{2m}$ (resp. $1\oplus{\sf MM}_{2m}$), then $(1\oplus W)g(\mathbf{Z})$ is an annihilator for ${\sf MM}_{2m+1}$ (resp. $1\oplus{\sf MM}_{2m+1}$). This shows the upper bound. Next we consider the lower bound. Suppose $g(W,\mathbf{Z})\neq 0$ is an annihilator for ${\sf MM}_{2m+1}(W,\mathbf{Z})$. We write $g(W,\mathbf{Z})$ as $g(W,\mathbf{Z})=Wg_{1}(\mathbf{Z})+g_{0}(\mathbf{Z})$. Noting that ${\sf MM}_{2m+1}(W,\mathbf{Z})=W\oplus{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$, we obtain
00 0 $=$ $g(W,\mathbf{Z}){\sf MM}_{2m+1}(W,\mathbf{Z})$
$=$ $g_{0}(\mathbf{Z}){\sf MM}_{2m}(\mathbf{Z})\oplus W\big{(}g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z})(1\oplus{\sf MM}_{2m}(\mathbf{Z}))\big{)}.$
So $g_{0}(\mathbf{Z}){\sf MM}_{2m}(\mathbf{Z})=0$ and $g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z})(1\oplus{\sf MM}_{2m}(\mathbf{Z}))=0.$ If $g_{0}$ is non-zero, then $g_{0}$ is an annihilator for ${\sf MM}_{2m}$ and so ${\sf deg}(g)\geq{\sf deg}(g_{0})\geq{\sf AI}({\sf MM}_{2m})$. If $g_{0}=0$, then since $g\neq 0$, it follows that $g_{1}\neq 0$. In this case, $g_{1}$ is an annihilator for $1\oplus{\sf MM}_{2m}$, and so ${\sf deg}(g)\geq 1+{\sf deg}(g_{1})\geq 1+{\sf AI}({\sf MM}_{2m})$. Consequently, in both cases ${\sf deg}(g)\geq{\sf AI}({\sf MM}_{2m})$.
On the other hand, if $g(W,\mathbf{Z})\neq 0$ is an annihilator for $1\oplus{\sf MM}_{2m+1}(W,\mathbf{Z})$, then noting that $W(1\oplus W)=0$, we obtain $g_{0}(\mathbf{Z})(1\oplus{\sf MM}_{2m}(\mathbf{Z}))=0$ and $g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z}){\sf MM}_{2m}(\mathbf{Z})=0$. If $g_{0}\neq 0$, then $g_{0}$ is an annihilator for $1\oplus{\sf MM}_{2m}$, and if $g_{0}=0$, then $g_{1}$ is an annihilator for ${\sf MM}_{2m}$. So again we have ${\sf deg}(g)\geq{\sf AI}({\sf MM}_{2m})$.
@@
We computed the algebraic immunity of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$ for small values of $n$ and observed that for $n\geq 4$, ${\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n})=1+\lfloor n/4\rfloor$. This suggests that the lower bound in Proposition 4 is exact if $n\not\equiv 0\bmod 4$ and if $n\equiv 0\bmod 4$, the actual algebraic immunity is one more than the lower bound. In a work [41] subsequent to the present paper, it was shown that for even $n$, ${\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n})\leq 1+2^{\lceil\log(n/4)\rceil}$. As mentioned in Section 2, for any Boolean function $f$, ${\sf FAI}(f)\geq 1+{\sf AI}(f)$. Our experiments suggest that for $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$, the fast algebraic immunity is actually one more than the algebraic immunity.
So $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$ does not have the best possible algebraic immunity. On the other hand, it can be implemented using $O(m)$ gates (as we show in Section 5). So it is possible to increase the number of variables to achieve the desired level of algebraic resistance without increasing the circuit size too much.
From Proposition 2 we have ${\sf MM}_{2m+1}$ is balanced, and the cryptographic resistance provided by ${\sf MM}_{2m+1}$ is very similar to that provided by ${\sf MM}_{2m}$. In particular, the linear bias of ${\sf MM}_{2m+1}$ is the same as that of ${\sf MM}_{2m}$, and the algebraic resistance of ${\sf MM}_{2m+1}$ is at least that of ${\sf MM}_{2m}$. In terms of implementation efficiency, ${\sf MM}_{2m+1}$ requires only one extra XOR gate in addition to the circuit to implement ${\sf MM}_{2m}$. We consider $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$, i.e. where $h$ is chosen to be ${\sf Maj}_{m}$ and $\pi$ is chosen to be ${\sf rev}$. The function ${\sf MM}_{2m+1}$ maps the all-zero string to 0. This can be a problem for use in a linear system, since such a system will map the all-zero string to the all-zero string. We modify ${\sf MM}_{2m+1}$ so that the all-zero string is mapped to 1. We define $f_{2m+1}:\{0,1\}^{2m+1}\rightarrow\{0,1\}$ to be

### v7

@@ 2 Preliminaries
By #S\#S we will denote the cardinality of a finite set $S$. The finite field of two elements will be denoted by $\mathbb{F}_{2}$, and for a positive integer $n$, $\mathbb{F}_{2}^{n}$ will denote the vector space of dimension $n$ over $\mathbb{F}_{2}$. By $\oplus$, we will denote the addition operator over both $\mathbb{F}_{2}$ and $\mathbb{F}_{2}^{n}$. An element of $\mathbb{F}_{2}^{n}$ will be considered to be an $n$-bit binary string.
For $n\geq 0$, let $\mathbf{x}=(x_{1},\ldots,x_{n})$ be an $n$-bit binary string. The support of $\mathbf{x}$ is ${\sf supp}(\mathbf{x})=\{1\leq i\leq n:x_{i}=1\}$, and the weight of $\mathbf{x}$ is ${\sf wt}(\mathbf{x})=\#{\sf supp}(\mathbf{x})$. By $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$ we will denote the all-zero and all-one strings of length $n$ respectively. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ be two $n$-bit strings. The distance between $\mathbf{x}$ and $\mathbf{y}$ is $d(\mathbf{x},\mathbf{y})=\#\{i:x_{i}\neq y_{i}\}$; the inner product of $\mathbf{x}$ and $\mathbf{y}$ is $\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}\oplus\cdots\oplus x_{n}y_{n}$; and we define $\mathbf{x}\leq\mathbf{y}$ if $x_{i}=1$ implies $y_{i}=1$ for $i=1,\ldots,n$.
@@
Let $f$ be an $n$-variable function. The algebraic normal form (ANF) representation of $f$ is the following: $f(X_{1},\ldots,X_{n})=\bigoplus_{\bm{\alpha}\in\mathbb{F}_{2}^{n}}a_{\bm{\alpha}}\mathbf{X}^{\bm{\alpha}}$, where $\mathbf{X}=(X_{1},\ldots,X_{n})$; for $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}_{2}^{n}$, $\mathbf{X}^{\bm{\alpha}}$ denotes the monomial $X_{1}^{\alpha_{1}}\cdots X_{n}^{\alpha_{n}}$; and $a_{\bm{\alpha}}\in\mathbb{F}_{2}$. If $f$ is not the zero function, then the (algebraic) degree of $f$ is ${\sf deg}(f)=\max\{{\sf wt}(\bm{\alpha}):a_{\bm{\alpha}}=1\}$. By convention, we assume that the degree of the zero function is 0. Functions of degree at most 1 are said to be affine functions. Affine functions with $a_{\mathbf{0}_{n}}=0$ are said to be linear functions.
@@
For two $n$-variable functions $f$ and $g$, the distance between them is $d(f,g)=\#\{\mathbf{x}\in\mathbb{F}_{2}^{n}:f(\mathbf{x})\neq g(\mathbf{x})\}$. The nonlinearity of an $n$-variable function $f$ is ${\sf nl}(f)=\min d(f,g)$, where the minimum is over all $n$-variable affine functions $g$. We define the linear bias of an $n$-variable Boolean function $f$ to be ${\sf LB}(f)=1/2-{\sf nl}(f)/2^{n}$.
@@
A function is said to be plateaued if its Walsh transform takes only the values $0,\pm v$, for non-zero $v$.
@@
Let $f$ be an $n$-variable function. The algebraic immunity of $f$ is defined [23, 44] as follows: ${\sf AI}(f)=\min_{g\neq 0}\{{\sf deg}(g):\mbox{ either }gf=0,\mbox{ or }g(f\oplus 1)=0\}$. From the restatement of Theorem 6.0.1 of [23] in the extended version of [23] available at http://ntcourtois.free.fr/toyolili.pdf (accessed on December 20, 2025), it follows that for any $n$-variable function $f$, there is a function $g$ of degree at most $\lceil n/2\rceil$ such that $h=fg$ is of degree at most $\lfloor n/2\rfloor$. From this result it easily follows that for any $n$-variable function $f$, ${\sf AI}(f)\leq\lceil n/2\rceil$. (One may observe that $fh=f^{2}g=fg$ and so $f(g\oplus h)=0$ which implies that $f$ has an annihilator of degree at most $\lceil n/2\rceil$).
The fast algebraic attack (FAA) was introduced in [22]. The idea of the attack is based on the following observation. Let $f$ be an $n$-variable function and suppose $g$ is another $n$-variable function of degree $e$ such that $gf$ has degree $d$. If both $e$ and $d$ are small, then $f$ is susceptible to an FAA. Given $f$, and for $e$ and $d$ satisfying $e+d\geq n$, it is known [22] that there exists functions $g$ and $h$ with ${\sf deg}(g)=e$ and ${\sf deg}(h)\leq d$ such that $gf=h$. For any pair of functions $g$ and $h$ of degrees $e$ and $d$ respectively, satisfying $gf=h$, we have $h=gf=gf^{2}=(gf)f=hf$ and so if $h\neq 0$, then $h$ is an annihilator of $1\oplus f$ which implies $d\geq{\sf AI}(f)$. If $1\leq e<{\sf AI}(f)$, then $g$ is not an annihilator of $f$ and so $h=gf\neq 0$, and we obtain $e+d\geq 1+{\sf AI}(f)$. The fast algebraic immunity (FAI) of $f$ is a combined measure of resistance to both algebraic and fast algebraic attacks: ${\sf FAI}(f)\allowbreak=\allowbreak\min\left(2{\sf AI}(f),\allowbreak\min_{g\neq 0}\allowbreak\{{\sf deg}(g)\allowbreak+\allowbreak{\sf deg}(fg):\allowbreak 1\allowbreak\leq\allowbreak{\sf deg}(g)\allowbreak<\allowbreak{\sf AI}(f)\}\right)$. For any function $f$, $1+{\sf AI}(f)\leq{\sf FAI}(f)\leq 2\,{\sf AI}(f)$.
@@ 3 Construction from Maiorana-McFarland Bent Functions
$(h,\pi)\mbox{-}{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$ $=$ $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})=\pi_{1}(\mathbf{X})Y_{1}\oplus\cdots\oplus\pi_{m}(\mathbf{X})Y_{m}\oplus h(\mathbf{X}).$ (2)
@@
${\sf AI}({\sf MM}_{2m})$ $\geq$ ${\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big(\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big).$ (3)
@@
$0$ $=$ $g(\mathbf{X},\mathbf{Y}){\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$ (4)
@@
Since the right hand side of (4) is equal to 0, for $\bm{\omega}\in\mathbb{F}_{2}^{m}$, the coefficient of $\mathbf{Y}^{\bm{\omega}}$ in the expansion on the right hand side of (4) must be equal to 0. Since $g(\mathbf{X},\mathbf{Y})\neq 0$, let $w\geq 0$ be the minimum integer such that there is an $\bm{\omega}^{\prime}$ with ${\sf wt}(\bm{\omega}^{\prime})=w$ and $g_{\bm{\omega}^{\prime}}(\mathbf{X})\neq 0$. In (4), equating the coefficient of $\mathbf{Y}^{\bm{\omega}^{\prime}}$ to 0, we have
$0$ $=$ $g_{\bm{\omega}^{\prime}}(\mathbf{X})\left(h(\mathbf{X})\oplus\left(\bigoplus_{i\in{\sf supp}(\bm{\omega}^{\prime}))}\pi_{i}(\mathbf{X})\right)\right)$
$=$ $g_{\bm{\omega}^{\prime}}(\mathbf{X})\big(\langle\bm{\omega}^{\prime},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big).$
So $g_{\bm{\omega}^{\prime}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\prime},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})$. Consequently, ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\prime})+{\sf deg}(g_{\bm{\omega}^{\prime}})\geq{\sf wt}(\bm{\omega}^{\prime})+{\sf AI}\big(\langle\bm{\omega}^{\prime},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big)$.
Now suppose that $g(\mathbf{X},\mathbf{Y})$ is an annihilator for $1\oplus{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$. Then again writing $g(\mathbf{X},\mathbf{Y})=\bigoplus_{\bm{\omega}\in\mathbb{F}_{2}^{m}}\mathbf{Y}^{\bm{\omega}}g_{\bm{\omega}}(\mathbf{X})$, an argument similar to the above shows that there is an $\omega^{\prime\prime}$ such that $g_{\bm{\omega}^{\prime\prime}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\prime\prime},\pi(\mathbf{X})\rangle\oplus 1\oplus h(\mathbf{X})$, and again we have ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\prime\prime})+{\sf AI}\big(\langle\bm{\omega}^{\prime\prime},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\big)$.
Combining the above two cases we get that there is a function $g$ which is either an annihilator of ${\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$, or an annihilator of $1\oplus{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$, and there is some $\bm{\omega}^{\star}\in\mathbb{F}_{2}^{n}$ such that ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big(\langle\bm{\omega}^{\star},\pi(\mathbf{X})\oplus h(\mathbf{X})\rangle\big)$. This shows (3).
@@
We consider a folklore extension of ${\sf MM}_{2m}$ to odd number of variables.
@@
The first point is well known. The second point is immediate from the definition of ${\sf MM}_{2m+1}$.
For the third point, we prove a more general statement than what is required. We show that if $f(\mathbf{Z})$ is an $n$-variable function and $h(W,\mathbf{Z})=W\oplus f(\mathbf{Z})$, then ${\sf AI}(f)\leq{\sf AI}(h)\leq 1+{\sf AI}(f)$.
Clearly if $g(\mathbf{Z})$ is an annihilator for $f(\mathbf{Z})$ (resp. $1\oplus f(\mathbf{Z})$), then $(1\oplus W)g(\mathbf{Z})$ is an annihilator for $f(\mathbf{Z})$ (resp. $1\oplus f(\mathbf{Z})$). This shows the upper bound. Next we consider the lower bound. Suppose $g(W,\mathbf{Z})\neq 0$ is an annihilator for $h(W,\mathbf{Z})$. We write $g(W,\mathbf{Z})$ as $g(W,\mathbf{Z})=Wg_{1}(\mathbf{Z})+g_{0}(\mathbf{Z})$. Noting that $h(W,\mathbf{Z})=W\oplus f(\mathbf{Z})$, we obtain
$0$ $=$ $g(W,\mathbf{Z})h(W,\mathbf{Z})$
$=$ $g_{0}(\mathbf{Z})f(\mathbf{Z})\oplus W\big(g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z})(1\oplus f(\mathbf{Z}))\big).$
So $g_{0}(\mathbf{Z})f(\mathbf{Z})=0$ and $g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z})(1\oplus f(\mathbf{Z}))=0.$ If $g_{0}$ is non-zero, then $g_{0}$ is an annihilator for $f$ and so ${\sf deg}(g)\geq{\sf deg}(g_{0})\geq{\sf AI}(f)$. If $g_{0}=0$, then since $g\neq 0$, it follows that $g_{1}\neq 0$. In this case, $g_{1}$ is an annihilator for $1\oplus f$, and so ${\sf deg}(g)=1+{\sf deg}(g_{1})\geq 1+{\sf AI}(f)$. Consequently, in both cases ${\sf deg}(g)\geq{\sf AI}(f)$.
On the other hand, if $g(W,\mathbf{Z})\neq 0$ is an annihilator for $1\oplus h(W,\mathbf{Z})$, then from $g(W,\mathbf{Z})(1\oplus h(W,\mathbf{Z}))=0$ and noting that $W(1\oplus W)=0$, we obtain $g_{0}(\mathbf{Z})(1\oplus f(\mathbf{Z}))=0$ and $g_{0}(\mathbf{Z})\oplus g_{1}(\mathbf{Z})f(\mathbf{Z})=0$. If $g_{0}\neq 0$, then $g_{0}$ is an annihilator for $1\oplus f$, and if $g_{0}=0$, then $g_{1}$ is an annihilator for $f$. So again we have ${\sf deg}(g)\geq{\sf AI}(f)$.
@@
The choice of $\pi$ as the bit reversal permutation instead of the identity permutation is to provide resistance to a type of differential attack which was proposed on a previous version. The attack is discussed in details in Section 4.4. See also Remark 1.
@@
We computed the algebraic immunity of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$ for values of $n\leq 16$ and observed that for $n\geq 4$, ${\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n})=1+\lfloor n/4\rfloor$. This suggests that the lower bound in Proposition 4 is exact if $n\not\equiv 0\bmod 4$ and if $n\equiv 0\bmod 4$, the actual algebraic immunity is one more than the lower bound. In a work [41] subsequent to the present paper, it was shown that for even $n$, ${\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n})\leq 1+2^{\lceil\log(n/4)\rceil}$. As mentioned in Section 2, for any Boolean function $f$, ${\sf FAI}(f)\geq 1+{\sf AI}(f)$. Our experiments suggest that for $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$, the fast algebraic immunity is actually one more than the algebraic immunity.
From the above discussion, it follows that the algebraic immunity of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{n}$ is less than the maximum possible value $\lceil n/2\rceil$. On the other hand, it can be implemented using $O(m)$ gates (as we show in Section 5). So it is possible to increase the number of variables to achieve the desired level of algebraic resistance without increasing the circuit size too much.
The function ${\sf MM}_{2m+1}$ is balanced (see Proposition 2), and the cryptographic resistance provided by ${\sf MM}_{2m+1}$ is very similar to that provided by ${\sf MM}_{2m}$. In particular, the linear bias of ${\sf MM}_{2m+1}$ is the same as that of ${\sf MM}_{2m}$, and the algebraic resistance of ${\sf MM}_{2m+1}$ is at least that of ${\sf MM}_{2m}$. In terms of implementation efficiency, ${\sf MM}_{2m+1}$ requires only one extra XOR gate in addition to the circuit to implement ${\sf MM}_{2m}$. We consider $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$, i.e. where $h$ is chosen to be ${\sf Maj}_{m}$ and $\pi$ is chosen to be $\sf rev$}. The function ${\sf MM}_{2m+1}$ maps the all-zero string to 0. This can be a problem for use in a linear system, since such a system will map the all-zero string to the all-zero string. We modify ${\sf MM}_{2m+1}$ so that the all-zero string is mapped to 1. We define $f_{2m+1}:\{0,1\}^{2m+1}\rightarrow\{0,1\}$ to be